OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)?

OpenStudy (binary3i):

y= mx+c m= (7-(-1))/(1-(-3)) m=2 y= 2x+c put the point to get c 1=2*7+c c=-13 so equation is y=2x-13 y intersept is when x=0 so y=-13

OpenStudy (anonymous):

Thank you so so much!

OpenStudy (binary3i):

you are always welcome